Okay I really need someone's help with this one. Substitution confuses me. I need to solve the system by substitution:
3x+3y=9
y=2x-12

Answer by checkley75(3666)

y=2x-12 now substitute (2x-12) for y in the other equation thus:
3x+3(2x-12)=9
3x+6x-36=9
9x=9+36
9x=45
x=45/9
x=5 answer.
y=2*5-12
y=10-12
y=-2 answer.
proof
3*5+3*-2=9
15-6=9
9=9